T/F if false make the statement right Every ratio tells us for every one of​ what’s...

T/F

if false make the statement right

  1. Every ratio tells us for every one of​ what’s on the​ top, here’s how many we have of​ what’s on the bottom.
  2. "Ratio", "proportion",​ "fraction" and​ "percent" all mean the same. thing.
  3. If the​ debt-equity ratio is​ 1.25, the equity multiplier would be 2.50.
  4. The higher the equity​ multiplier, the greater is the proportion of a​ firm’s assets that are financed with equity.
  5. Common-size values on the balance sheet show each item as a percent of total equity.
  6.  If sales increase by​ 5% and total assets fall by​ 1%, the TAT ratio would go down by approximately​ 6%
  7.  A decrease in the current ratio indicates an improvement in a​ firm’s liquidity.
  8.  An increase in the cash coverage ratio means that a firm is less likely to default on its outstanding debt.
  9.  Ceteris paribus​, according to the DuPont​ framework, an increase in the use of debt would increase a​ firm’s ROE.
  10.  For firms with lower​ P/E ratios, investors are valuing each dollar of earnings less than for firms with higher​ P/E ratios.
  11.  Corporate managers who are doing a better job of serving owners would see the​ market-book ratio their firm exceed the ratio for managers who are not doing as good a job.
  12. Ceteris paribus​, the cash coverage ratio and the amount of debt are directly related.
Answer #1

1.
"Ratio", "proportion",​ "fraction" and​ "percent" all mean the same. thing.: TRUE

2.
If the​ debt-equity ratio is​ 1.25, the equity multiplier would be 2.50. FALSE (Equity multiplier=(Debt/Equity+1)/1)=2.25

3.
The higher the equity​ multiplier, the greater is the proportion of a​ firm’s assets that are financed with equity. FALSE (Debt and not equity)

4.
Common-size values on the balance sheet show each item as a percent of total equity. FALSE (as a percent of total assets and not total equity)

5.
If sales increase by​ 5% and total assets fall by​ 1%, the TAT ratio would go down by approximately​ 6% FALSE (=Sales/Total Assets=1.05/0.99=1.06) (up and not down)

6.
A decrease in the current ratio indicates an improvement in a​ firm’s liquidity. FALSE (deterioration and not improvement)

7.
An increase in the cash coverage ratio means that a firm is less likely to default on its outstanding debt. FALSE (more likely)

8.
Ceteris paribus​, according to the DuPont​ framework, an increase in the use of debt would increase a​ firm’s ROE. TRUE

9.
For firms with lower​ P/E ratios, investors are valuing each dollar of earnings less than for firms with higher​ P/E ratios. TRUE

10.
Corporate managers who are doing a better job of serving owners would see the​ market-book ratio their firm exceed the ratio for managers who are not doing as good a job. TRUE

11.
Ceteris paribus​, the cash coverage ratio and the amount of debt are directly related. FALSE (inversely)